# The effects of fibre spatial distribution and relative orientation on the percolation and mechanics of stochastic fibre networks: A model of peptide hydrogels

### v1

@@ Abstract
We first studied percolation. All the subsequent analysis was carried out on percolated systems. Network percolation was investigated based on the work of Li and Zhang [11], who showed that, $\rho_{0.5}(L$) (the density at which the percolation probability is 0.5 for finite $L$) converges to $\rho_{c}$ as $L\rightarrow\infty$, since
@@
where $\rho_{c}$, $\nu_{c}$, $b_{0}$, and $a_{1}$ are the critical percolation density, the correlation-length exponent (a universal constant that is only affected by dimensionality, which equals $4/3$ for 2D systems [11, 22]), and some fitting constants, respectively (see Supplemental Material for further details).
The percolation thresholds are presented in Table 1. As we modify the spatial dispersion, or make the fibres more likely to be perpendicular, $\rho_{c}$ decreases. If we make the fibres more likely to be parallel, $\rho_{c}$ increases.

### v2

@@ Abstract
We first studied percolation. All the subsequent analysis was carried out on percolated systems. Network percolation was investigated based on the finding that $\rho_{0.5}(L)$ (the density at which the percolation probability is 0.5 for finite $L$) converges to $\rho_{c}$ as $L\rightarrow\infty$ [6, 24, 25], since
@@
where $\rho_{c}$, $\nu_{c}$, $b_{0}$, and $a_{1}$ are the critical percolation density, the correlation-length exponent (a universal constant that is only affected by dimensionality, which equals $4/3$ for 2D systems [6, 8]), and some fitting constants, respectively. We used a method similar to that developed by Pike and Seager [26] to find $\rho_{0.5}(L)$ (see Supplemental Material for further details).
The percolation thresholds are presented in Table 1. Our results for Mikado networks are consistent with the literature [6, 27]; however, as we modify the spatial dispersion, or make the fibres more likely to be perpendicular, $\rho_{c}$ decreases, and if we make the fibres more likely to be parallel, $\rho_{c}$ increases.